4. During five years, the inflation rates are 3.8%, 9.2%, 2.1%, 7.4%, 5.3%. Calculate the average annual inflation rate. (Hint: A question like this one always requires taking a geometric mean of the (1+ R) values, to get (1+R,).) 5. In a sample of N 100 observations, D, is a dummy variable that takes a value of 1 for 64% of the sample, and 0 for the other 36%. Calculate the mean and standard deviation in this variable.
Answer #1

Question 4

here the average annual inflation rate would be calculated by using the given formula

1 + \bar R_g = [(1 + 0.038) * (1 + 0.092) * (1 + 0.021) * (1 + 0.074) * (1 + 0.053)]1/5

1 + \bar R_g = 1.0553

\bar R_g = 0.0553 or 5.53%

QUeestion 5

Here N = 100

Here

the mean = (1 * 64 + 0 * 36)/100 = 0.64

Variance = 1/100[64 * (1 - 0.64)2 + 36 * (0 - 0.64)2] = 0.2304

standard deviation = sqrt(0.2304) = 0.48